# Agnostics: Learning to Code in Any Programming Language via Reinforcement with a Universal Learning Environment

### v1

@@ 4 Evaluation
To evaluate our approach and codebase, we train and benchmark multiple models on 5 low-resource programming languages. We measure pass@1 accuracy with reasoning disabled and 20 samples per prompt at temperature 0.2.
@@ 4.1 Training datasets
Ag-Codeforces-X, the main dataset we use for training, was created based on competitive programming problems from the Open-R1 Codeforces dataset (Penedo et al. 2025). Few adjustments were necessary, since the problems already specified programs and tests using standard I/O. The train split contains 5369 problems. See § B for more details.
@@ 4.2 Benchmarks
MultiPL-E (Cassano et al. 2023) is a well-established benchmark which has been frequently used to evaluate the performance of new LLMs on a broad set of languages (e.g., Kimi Team (2025); Yang et al. (2025); Grattafiori et al. (2024); ByteDance et al. (2025)). MultiPL-E was prepared by compiling HumanEval (Chen et al. 2021) prompts and unit tests from Python to each target language. Each MultiPL-E programming language requires a $\approx 500$ LOC prompt and test translator, considerably more effort than writing an Agnostics configuration file.
A major limitation of MultiPL-E is being too easy for frontier models. With Python, frontier models are now evaluated on solving programming contest problems (Jain et al. 2024a); no multi-language benchmarks are as challenging.
Ag-LiveCodeBench-X, a contribution of this paper, is a new multi-language benchmark derived from LiveCodeBench. LiveCodeBench 5.0 has 880 problems, of which 381 have Python starter code and test cases. The remaining 499 problems do not use starter code and instead use standard I/O to specify and test solutions. Hence we used these problems to transform LiveCodeBench into an Agnostics dataset. Accordingly, benchmarking a new programming language with Ag-LiveCodeBench-X is straightforward: we can reuse the language configurations and execution environment from our trainer (§ 3.3).
Importantly, the problems from LiveCodeBench and thus also Ag-LiveCodeBench-X are filtered out from our training sets. Moreover, our results show that Ag-LiveCodeBench-X is significantly harder than MultiPL-E.
@@ 4.3 Results
We now present our results. We use a few abbreviations in the tables. Ag-LCB-X stands for Ag-LiveCodeBench-X; we clarify abbreviated model names in the text. One table row may present scores of many models; the column for language X shows the score of the model trained on X.
Table 1: Pass@1 rates on AG-LiveCodeBench-X for Qwen 3 4B models trained on five low-resource languages using Ag-Codeforces-X. Each model achieves a SOTA result for an open-weight model with ${\leq}16$B parameters.
Model Ag-LiveCodeBench-X
X= Lua Julia R OCaml Fortran
Llama-3.3-70B-Ins 25 22 13 7 3
Qwen3-32B 22 26 17 2 1
DSCv2-Lite-Ins-16B 13 12 9 7 6
@@
Table 2: Pass@1 rates on MultiPL-E for Qwen 3 4B models trained using Ag-Codeforces-X. The significant improvement indicates that they did not overfit to the competitive programming format.
@@
Qwen3-4B 61 51 36
[…]
SOTA small LLMs for low-resource PLs
Using Agnostics, we train Qwen 3 4B on Ag-Codeforces-X specialized to Fortran, Julia, Lua, OCaml, and R. The resulting Qwen3-4B-CF-X models are, to the best of our knowledge, state-of-the art low-resource programming language code models with open-weights and ${\leq}16$B parameters.
Benchmarking them on Ag-LiveCodeBench-X (table 1), we see significant improvements. (i) On every language, the models match or outperform DeepSeek Coder v2 Lite Instruct (16B). (ii) On every language, our performance comes close to or even exceeds that of Qwen 3 32B and Llama 3.2 70B. (iii) Compared to the base model, Qwen 3 4B, pass@1 improves by a factor of 1.5–2x. It is safe to assume that the Qwen models, like the Llama models (Grattafiori et al. 2024), are trained on all the publicly available coding data; hence, Agnostics improves the models beyond what typical training on such data allows.
(iv) Finally, on OCaml and Fortran, we improve the performance of Qwen 3 4B from near zero to 7% and 15%. Notably, although we train the models with a prompt prefix to facilitate learning (§ 3.1), we do not use this prompt prefix for evaluation. Thus the pass@1 scores represent what the model learned, and not information provided in context.
A potential weakness of our approach is that our models are overspecialized to the competitive programming format, and can only generate programs that use standard I/O. We also evaluate with the established MultiPL-E benchmark, which does not use this format, and also find significant improvements (table 2). [^2] Note that MultiPL-E does not support OCaml and Fortran. This indicates that our models have not overfit to the competition programming format.
Figure 5: Training Qwen3-4B-CF-X, GRPO batch pass@1.
Figure 5 shows the GRPO batch pass@1 rates seen when training Qwen3-4B-CF-X. All the models follow similar curves, partially due to being trained on the same data permutation. Nearly all the models slowly keep improving almost until the dataset end. We also observed the train and test split rewards to be correlated with each other (§ D.2).
Table 3: Pass@1 rates for larger Qwen 3 8B models trained on three languages using Ag-Codeforces-X. Agnostics scales to models of this size.
Model MultiPL-E Ag-LCB-X
X= Lua Julia R X= Lua Julia R
Qwen3-8B 63 53 44 11 9 9
[…]
Agnostics scales to larger models
To test if improvements from Agnostics training scale with model size, we train the Qwen 3 8B model on Ag-Codeforces-X specialized to Lua, Julia and R and benchmarked it on Ag-LiveCodeBench-X and MultiPL-E (Table 3). The Qwen3-8B-CF-X models improve significantly on both benchmarks, improving over their 4B counterparts.
We expect Agnostics to scale to even larger models, with appropriate computing resources. However, we found that Agnostics training with Ag-Codeforces-X does not improve two smaller models: Qwen 3 1.7B and Llama 3.2 3B Instruct, perhaps due to the problems being too difficult.
Table 4: Pass@1 rates on for Qwen 3 4B models trained using MBPP. This smaller and easier dataset still results in improvements, though smaller than those in Tables 1 and 2.
Model MultiPL-E Ag-LCB-X
X= Lua Julia R X= Lua Julia R
Qwen3-4B 51 61 36 11 10 10
[…]
Agnostics works with easier problems
So far, all of the models that we have trained so far have used the Open-R1 Codeforces dataset. To show that Agnostics works with other datasets, we also train models for Julia, Lua, and R using the MBPP training set. (§ 3.1 describes how we prepare MBPP.) Note that the MBPP problems are trivial—see figure 3(a)—compared to the Open-R1 Codeforces problems. So, we cannot expect these models trained on the MBPP problems to be as good as the models we presented before.
Nevertheless, training on MBPP still improves performance on Lua and Julia (table 4). On R, the table shows a small drop in performance on Ag-LiveCodeBench-X, but an improvement on MultiPL-E.
Table 5: Pass@1 rates for DeepSeek Coder V1 and Phi 4 models trained using Ag-Codeforces-X. We find that Agnostics works on other model families.
Model MultiPL-E Ag-LCB-X
X= Lua Julia R X= Lua Julia R
Phi4-mini-ins 40 39 34 7 7 0
DSC-6.7B-Ins 40 54 37 8 5 7
@@
Agnostics works on other model families
@@
Finally, to test if Agnostics works on other models, we train both DeepSeek Coder 6.7B Instruct (Guo et al. 2024) and Phi 4 Mini Instruct (Microsoft et al. 2025) with Ag-Codeforces-X specialized to Julia, Lua, and R. Agnostics is able to improve the performance of both models on all languages, as measured by MultiPL-E and Ag-LiveCodeBench-X with one exception (Table 5). Phi 4 Mini Instruct scores zero on Ag-LiveCodeBench-R, and training does not improve the score. However, on the easier MultiPL-E benchmark, we see that the model is slightly better on simpler R problems ($34\rightarrow 35$).
@@
Note that DeepSeek Coder 6.7B is a relative old LLM, superseded by the much larger DeepSeekV2 and V3 models. Unlike Qwen 3, DeepSeek Coder is not trained with reinforcement learning, but is only an instruction-tuned model. Thus this result also shows that Agnostics can work on models that have had relatively limited post-training.
@@
Figure 6: We classify with an LLM what bugs appear in programs produced by Qwen 3 4B and our trained model, Qwen3-4B-CF-OCaml. Each radial represents a bug category, and points along the radial indicate how many programs appear to have that bug. Qwen 3 4B makes more fundamental programming mistakes, such as syntax errors and misusing builtins, indicating its limited grasp of OCaml. Our trained model makes significantly fewer such mistakes. We see a small increase in logic flaws: the trained model makes fewer shallow mistakes, revealing deeper issues.
@@ 4.4 Qualitative Improvements
In this section we take a deeper look at how training with Agnostics addresses the kinds of errors that Qwen 3 4B makes on low-resource languages. First, we define a taxonomy of common bugs by prompting an LLM (o3) to classify bugs in a sample of faulty programs and then lightly editing the suggestion. § E has the full taxonomy and the prompt we used to develop it. The taxonomy spans fundamental programming errors, such as syntax errors, and more subtle mistakes such as logic flaws.
We then sample 100 problems from Ag-LiveCodeBench-X, and take five programs produced by Qwen 3 4B and the Agnostics models trained on Open-R1 Codeforces problems. Recall from table 1 that these are significantly better than Qwen 3 4B. Using Sonnet 4, we use the taxonomy to classify the bugs in each program.
Figure 6 shows the bug distribution for OCaml. We see that the base Qwen 3 4B models makes substantially more fundamental programming mistakes in OCaml. More programs have syntax errors (55% vs 35% after training), more programs misuse builtin functions (60% vs 32%), and so on. However, we observe a small increase in logic flaws (18% vs 25% after training). Inspecting the results, the reason for this is that when a program is full of syntax errors and hallucinated functions, it is very difficult to even determine whether or not the algorithmic approach is correct. Training eliminates these shallow bugs and lets deeper issues manifest. We see the same patterns across the other four languages (§ E).
@@ 5 Conclusion
LLMs are strongest where pre-training and post-training data are abundant, and weakest where practitioners arguably need them most: for low-resource programming languages. This paper proposes Agnostics, a language-agnostic post-training pipeline that removes the per-language engineering tax by verifying code purely via externally observable behavior. A single verifier and a short configuration are enough to adapt the same reinforcement learning setup to new languages.
@@
Empirically, Agnostics consistently improves small open-weight models on five low-resource languages–Lua, Julia, R, OCaml, and Fortran–without requiring language-specific test translators. Training Qwen-3 4B with Ag-Codeforces-X yields large gains on our new Ag-LiveCodeBench-X benchmark and on MultiPL-E, often rivaling or surpassing 16B–70B open-weight baselines. The method scales to larger and different model families: Qwen-3 8B benefits similarly, and we observe improvements on DeepSeek Coder 6.7B and Phi-4 Mini as well. Error-type analysis shows that the gains come from reducing fundamental language-usage mistakes.
@@
A practical advantage of the approach is how little per-language work it requires. After the framework was in place, adding OCaml and Fortran took us less than an hour each. We expect adoption to be just as straightforward for any language with a command-line toolchain.
@@
We see no reason to expect the approach to not work for models larger than 8B, although our experiments are limited by available compute. For scaling data, the Agnostics reformulation strategy apply to much larger problem sets as well. For instance, OpenCodeReasoning has $\sim$600K problems with Python solutions (Ahmad et al. 2025); converting such corpora into language-agnostic I/O tasks would provide rich RL datasets with many target languages with minimal additional engineering.
6 Acknowledgments

### v2

@@ 4 Evaluation
To evaluate Agnostics, we train and benchmark models on 5 low-resource programming languages. We measure pass@1 accuracy with reasoning disabled, 20 samples per prompt at temperature 0.2. We trained the models for 1 epoch, unless specified otherwise.
@@ 4.1 Training datasets
Ag-Codeforces-X, the main dataset we use for training, was created based on competitive programming problems from the Open-R1 Codeforces dataset (Penedo et al., 2025). Few adjustments were necessary, since the problems already specified programs and tests using standard I/O. The train split contains 5369 problems.
Ag-MBPP-X, the other training dataset we use, was created from MBPP as explained in § 3.1.
Both datasets were analyzed to ensure no contamination with benchmarks we use. See § A for a more detailed description of how the datasets were prepared.
@@ 4.2 Benchmarks
MultiPL-E (Cassano et al., 2023) is a well-established benchmark, frequently used to evaluate the performance of new LLMs on a broad set of languages (e.g., Kimi Team (2025); Yang et al. (2025); Grattafiori et al. (2024); ByteDance et al. (2025)). MultiPL-E was prepared by compiling HumanEval (Chen et al., 2021) prompts and unit tests from Python to each target language. Each MultiPL-E programming language requires a ${\approx}500$ LOC prompt and test translator, considerably more effort than writing an Agnostics configuration file. A major limitation of MultiPL-E is being too easy for frontier models. With Python, frontier models are now evaluated on solving programming contest problems (Jain et al., 2024b); no multi-language benchmarks are as challenging.
Ag-LiveCodeBench-X, a contribution of this paper, is a new multi-language benchmark derived from LiveCodeBench. (The benchmark has no overlap with our training data; see § A.) LiveCodeBench 5.0 has 880 problems; 381 problems have Python starter code and test cases, while the remaining 499 problems instead use standard I/O to specify and test solutions. We used the latter problems to transform LiveCodeBench into an Agnostics dataset. Accordingly, benchmarking a new programming language with Ag-LiveCodeBench-X is straightforward: we can reuse our trainer’s language configurations and execution environment (§ 3.3). Our results and leaderboard (see § 1) show that Ag-LiveCodeBench-X is far harder than MultiPL-E and is challenging for frontier models as of February 2026.
@@ 4.3 Results
Model Ag-LCB-X
X= Lua Julia R
Llama 3.3 70B Ins 25 22 13
Qwen 3 32B 22 26 17
DSC v2 Lite Ins 16B 13 12 9
[…]
Qwen 3 8B 11 9 9
@@
[…]
Table 1: Ag-LCB-X pass@1.
We now present our results. We use a few abbreviations in the tables. Ag-LCB-X stands for Ag-LiveCodeBench-X; we clarify abbreviated model names in the text. Highlighted rows present our models; note that each cell in such a row presents the score of a different model trained on programming language X. We compute the score as explained in § 4.
SOTA small LLMs for low-resource PLs
Using Agnostics, we train Qwen 3 4B and 8B on Ag-Codeforces-X specialized to Fortran, Julia, R, Lua, and OCaml. The resulting models are state-of-the-art at coding in low-resource programming languages among open-weight models with ${\leq}16$B parameters.
Model Ag-LCB-X
X= OCaml Fortran
Sonnet 4 6 6
Llama 3.3 70B Ins 7 3
Qwen 3 32B 2 1
DSC v2 Lite Ins 16B 7 6
[…]
Qwen 3 8B 0 0
[…]
Table 2: Ag-LCB-X pass@1.
Agnostics training yields significant improvements on Ag-LiveCodeBench-X (tables 1 and 2). (i) On every language, our models match or outperform DeepSeek Coder v2 Lite Instruct (16B), and their performance comes close to or exceeds that of Qwen 3 32B and Llama 3.2 70B. (ii) The OCaml and Fortran scores improve from near zero to ${\approx}7\%$ and ${\approx}15\%$, outperforming even some frontier models. Importantly, these scores show what the models learned from training: during evaluation we omit the prompt prefix used to facilitate learning (§ 3.1). (iii) Finally, the pass@1 scores improve by a factor of 1.5–2x over the base model. This indicates that Agnostics can improve models beyond usual training on all the publicly available code data, as we can safely assume that the Qwen models were trained on all such data, like the Llama models (Grattafiori et al., 2024).
@@
Qwen 3 4B 61 51 36
Qwen 3 8B 63 53 44
@@
[…]
Table 3: MultiPL-E pass@1.
Models trained with our approach generalize over the competitive programming format: the improvements are not limited to synthesizing programs using standard I/O. To demonstrate this, we evaluate them on the established MultiPL-E benchmark. It features problems which ask for Python functions operating on usual Python data structures, and we find that our training also significantly improves the models on such problems (table 3). [^2] Note that MultiPL-E does not support OCaml and Fortran. We also confirmed that our training does not lower performance on other programming languages (§ C.4).
Figure 5 shows the GRPO batch pass@1 rates seen when training Qwen3-4B-CF-X. All the models follow similar curves, partially due to being trained on the same data permutation. Nearly all the models slowly keep improving almost until the dataset end. We also observed the train and test split rewards to be correlated with each other (§ C.2).
Figure 5: Group batch pass@1, Qwen3-4B-CF-X.
Agnostics scales to larger models
To test if the gains from Agnostics training scale with model size, we train the Qwen 3 8B model on Ag-Codeforces-X specialized to Lua, Julia and R and benchmarked it on Ag-LiveCodeBench-X and MultiPL-E (tables 1 and 3). The Qwen3-8B-CF-X models show significant gains on both benchmarks, improving over their 4B counterparts. We expect Agnostics to scale to even larger models, with appropriate computing resources. However, we found that Agnostics training on Ag-Codeforces-X does not improve two smaller models, Qwen 3 1.7B and Llama 3.2 3B Instruct, perhaps due to the problems being too difficult for models of this size.
Agnostics works with easier problems
All of the models we discussed so far were trained on Ag-Codeforces-X. To show that Agnostics works with other datasets, we also train models for Julia, Lua, and R using the MBPP training set. (§ 3.1 describes how we prepare MBPP.) MBPP problems are trivial compared to the Open-R1 Codeforces problems (see figure 3(a)), and we cannot expect models trained on the MBPP problems to be as good as ones we presented before. Still, training on MBPP improves Lua and Julia performance (tables 1 and 3). The table shows a small drop in R performance on Ag-LiveCodeBench-X, but a significant improvement on MultiPL-E.
Model MultiPL-E Ag-LCB-X
X= Lua Julia R Lua Julia R
SmolLM3 3B 11 12 18 1 2 2
Phi 4 mini ins 40 39 34 8 8 5
DSC 6.7B Ins 40 54 37 8 5 8
[…]
Table 4: Non-Qwen models, pass@1.
Agnostics works on multiple model families
@@
To show that Agnostics works on non-Qwen models, we train SmolLM 3 (SmolLM3 Team, 2025), Phi 4 Mini Instruct (Microsoft, 2025), and DeepSeek Coder 6.7B Instruct (Guo et al., 2024) on Ag-Codeforces-X specialized to Lua, Julia, and R. Agnostics improves these models’ performance on all languages, as measured by MultiPL-E and Ag-LiveCodeBench-X (Table 4).
@@
Note that DeepSeek Coder 6.7B is a relatively old LLM, superseded by the much larger DeepSeekV2 and V3 models. Unlike Qwen 3, DeepSeek Coder is not trained with reinforcement learning, but is only an instruction-tuned model. Thus this result also shows that Agnostics can work on models that have had relatively limited post-training.
Model Ag-LCB-Fortran
Sonnet 4 Thinking (teacher) 12
Qwen 3 4B (student) 0
[…]
Table 5: Distillation experiment results.
Agnostics outperforms distillation
So far we discussed training a model on its generations. An alternative is to distill a larger model (assuming one exists). As larger models do not perform very well on many low-resource programming languages, one can expect distillation to be less effective.
We run the following experiment to verify this claim. Using Sonnet 4 Thinking, we synthesize Fortran solutions to Ag-Codeforces-X problems, creating a training set of 1,987 items. (For 13 items, Sonnet 4 (sonnet-4-20250514) with extended thinking does not produce a response within its reasoning budget.) To make sure generating the training items does not use significantly more compute compared to Agnostics training, we use at most 32K reasoning tokens, spending $\mathdollar 96$ to generate the items. We fine-tune Qwen 3 4B for 3 epochs (batch size 64, learning rate $2\times 10^{-5}$, cosine learning rate decay with warmup ratio $0.1$). Table 5 shows the resulting models reach scores far lower than the $15\%$ of Qwen3-4B-CF-Fortran (table 2).
Agnostics outperforms rejection sampling
We also ran a small experiment to confirm that rejection sampling would be prohibitively expensive in our case. In rejection sampling with supervised fine-tuning , we prompt the model to synthesize $n$ solutions to each task, reject solutions that fail tests, and fine-tune on the task-solution pairs that pass tests. Cassano et al. (2024) use this approach to get solutions for ${\approx}30\%$ of the tasks in their dataset.
The efficiency of this approach depends on the hardness of the task and the capabilities of the model. In this paper, we use newer models that are marginally better at low-resource languages (based on MultiPL-E benchmark results). The task of Cassano et al. (2024) is to translate a simple, self-contained Python function from the model’s pretraining data into an equivalent function in another programming language. This task is significantly easier than the Ag-Codeforces-X task, which is to solve a competitive programming problem in a low-resource language without any reference code.
Rejection sampling would be prohibitively expensive for the low-resource programming languages we consider. Cassano et al. (2024) report a $30\%$ success rate on their code translation task. During Agnostics training, Qwen3-4B-CF-Fortran generated a correct answer to a train split problem only $6.64\%$ of the time, generating $11400$ verified programs overall. We also sampled responses to the same problems from the base model of Qwen3-4B-CF-Fortran, Qwen 3 4B, taking the same amount of samples with the same generation parameters as used during training. The base model succeeded $0.09\%$ of the time, generating only $158$ test-passing programs.
@@
Figure 6: LLM labels of bugs in programs synthesized by Qwen 3 4B and our trained model, Qwen3-4B-CF-OCaml. A radial represents a bug class, and points along the radial show how many programs seem to have that bug. Qwen 3 4B makes more fundamental programming mistakes, such as syntax errors and misusing builtins, showing its limited grasp of OCaml. Our trained model makes far fewer such mistakes. We see a small increase in logic flaws: the trained model makes fewer shallow mistakes, revealing deeper issues.
@@ 4.4 Qualitative Improvements
We now take a deeper look at how training with Agnostics qualitatively improves Qwen 3 4B. First, we define a taxonomy of common bugs; we prompt o3 to classify bugs in a sample of faulty programs and lightly edit the suggestion. § D has the full taxonomy and the prompt used to develop it. The taxonomy spans fundamental programming errors such as syntax errors, and subtler mistakes such as logic flaws.
We sample 100 problems from Ag-LiveCodeBench-X, and for each take five OCaml programs produced by Qwen 3 4B and the models trained on Ag-Codeforces-X. Recall from table 2 that our models significantly outperform the base Qwen 3 models on our benchmark. With Sonnet 4, we use the taxonomy to classify the bugs in each program.
Figure 6 shows the bug distribution for OCaml. We see that the base Qwen 3 4B model makes many more fundamental OCaml programming mistakes. More programs have syntax errors (55% vs 35% after training), more programs misuse builtin functions (60% vs 32%), and so on. We do observe a small increase in logic flaws (18% vs 25% after training). When a program is full of syntax errors and hallucinated functions, it is hard to decide if the algorithmic approach is correct; training eliminates these shallow bugs and lets deeper issues manifest. Models trained for the other four programming languages show the same patterns (§ D).
@@ 5 Conclusion
Arguably, practicioners need LLMs the most for low-resource programming languages, which need highly specialized knowledge. Unfortunately, this is precisely where LLMs are weak due to a lack of sufficient pre-training and post-training data. We propose Agnostics, a language-agnostic post-training pipeline that minimizes the per-language engineering tax by verifying code entirely via externally observable behavior. A short configuration file is enough to adapt the pipeline, including the reinforcement learning setup, to a new programming language.
@@
Empirically, Agnostics consistently improves small open-weight models on five low-resource languages—Lua, Julia, R, OCaml, and Fortran—without requiring language-specific test translators. Training Qwen 3 4B with Ag-Codeforces-X yields large gains on our new Ag-LiveCodeBench-X benchmark and on MultiPL-E, often rivaling or surpassing 16B–70B open-weight baselines. The method scales to larger and different model families: Qwen 3 8B shows similar gains to its smaller sibling, and we also observe improvements on DeepSeek Coder 6.7B, Phi 4 Mini and SmolLM3 3B. Error-type analysis shows our training decreases fundamental programming language mistakes.
@@
A practical advantage of the approach is how little per-language work it requires. After the framework was in place, extending the pipeline to support OCaml and Fortran took us less than an hour each. We expect adaption to be just as straightforward for any pragmatic programming language with a command-line toolchain.
@@
We believe the approach scales to models of arbitrary size, although our experiments are limited by available compute to at most 8B models. For scaling data, the Agnostics reformulation approach also applies to much larger problem sets. For instance, OpenCodeReasoning has $\sim$600K problems with Python solutions (Ahmad et al., 2025); converting such corpora into language-agnostic I/O tasks would provide rich RL datasets with many target languages with minimal additional engineering.
Reproducibility Statement
We provide references to our data, code, and artifacts at agnostics.abgru.me. We share our datasets and models at huggingface.co/collections/nuprl/agnostics, our Wandb training logs at wandb.ai/nuprl/Agnostics, and our training framework at github.com/nuprl/agnostics-framework.
The existing datasets we used are publicly available and are accompanied by citations. We publicly release all datasets we introduce, allowing free use for research. In the same way, we release all code required for conducting and analyzing our experiments, including the code for dataset preparation, as well as the models we presented and Wandb records from training them. We state the number and range of values tried per (hyper) parameter, and outline how we chose the final values and what they are (§§ C.1 and 3.3). We specify the computing infrastructure (hardware and software) we used for our experiments (§ C.5). The released codebases specify the exact versions of all the libraries we used.
Acknowledgments